Does there exist a plane targent to x²-2y²-2z²=8 and which passes through
2x+3y+2z = 8, x−y+2z = 5? Justify your answer.
1
Expert's answer
2020-03-06T10:53:17-0500

The equation of the straight line is

"2x+3y+2z=8\\
x-y+2z=5"

the direction vector

"\vec{a}=[\vec{n_1},\vec{n_2}]\\
\vec{n_1}=(2;3;2)\\
\vec{n_2}=(1;-1;2)\\
\vec{a}=\left(
\begin{vmatrix}
 3 & 2 \\
 -1 & 2
\end{vmatrix};-\begin{vmatrix}
 2 & 2 \\
 1 & 2
\end{vmatrix};\begin{vmatrix}
 2 & 3 \\
 1 & -1
\end{vmatrix}\right)=\\
=(8;-2-5)"

 point

let "y=0" , then

"2x+2z=8\\
x+2z=5\\
x=3\\
z=1"

"A(3,0,1)"

The equation of a straight line is

"\frac{x-3}{8}=\frac{y-0}{-2}=\frac{z-1}{-5}"

Let's take an arbitrary point "M(x_0;y_0;z_0)" on the surface

"x^2-2y^2-2z^2=8\\
x_0^2-2y_0^2-2z_0^2=8\\"

 The equation of a plane is

"\begin{vmatrix}
 x-x_0 & y-y_0&z-z_0 \\
 3-x_0 & 0-y_0& 1-z_0\\
8&-2&-5
\end{vmatrix}=0"  

"(x-x_0)(5y_0+2-2z_0)+(y-y_0)(15-5x_0+8-8z_0)+\\
+(z-z_0)(-6+2x_0+8y_0)=0\\
(x-x_0)(5y_0-2z_0+2)+(y-y_0)(-5x_0-8z_0+23)+\\
+(z-z_0)(2x_0+8y_0-6)=0\\"

hence the plane exists.